Help please? I need to know how to find line in point slope form for:
x intercept =-1/2
y intercept = 3
Thanks,
Anj

Answer by Alan3354(69443)

I need to know how to find line in point slope form for:
x intercept =-1/2
y intercept = 3
--------------
If x-int = (A,0) and y-int = (0,B), then
Bx + Ay = A*B
-------------
You might want to remember that, or not. There's no end of things you can memorize.
----
3x - y/2 = -3/2
6x - y = -3 (standard form)
------
Or,
y = 6x + 3 (slope-intercept form)
=============
(Not sure what point slope form is)